# LayerIF: Estimating Layer Quality for Large Language Models using Influence Functions

### v1

@@ 1 Introduction
There is a broad effort in the ML community to interpret and explain the behavior of these complex models. Many approaches take a model-centric perspective, examining the learned weights, activations, or attention patterns [20, 21, 22, 23] within the network to deduce how it works. In contrast to these methods that probe the internals, Influence Functions (IFs) [24, 25] offer a data-centric lens on model behavior. Rather than focusing on internal representations, Influence Functions (IFs) measure how a model’s output would change if a given training example was perturbed or removed, effectively tracing the influence of training data on the model’s predictions. They have proven effective in a range of downstream machine learning tasks, including mislabeled data detection [26, 27, 28], optimal subset selection [29, 30, 31], model interpretation [32, 33, 34], data attribution [35], data valuation [36], in-context learning [37], and analyzing model biases [38, 39].
@@
Figure 1: We present an overview of the LayerIF pipeline via the green arrows. We quantify per-layer quality in a pretrained LLM using Influence Functions (IFs), demonstrating that the same model produces distinct layer-wise IF scores across different datasets, revealing dataset-specific specialization. The pipeline begins by extracting gradients at each Transformer block of the LLM, computed separately for each dataset. These gradients and the dataset are then used to estimate layer-wise IF scores that serve as data-driven proxies for layer quality. To demonstrate the utility of these scores, we consider two downstream applications: (a) the allocation of optimal experts per layer in a LoRA-MoE architecture, and (b) the computation of structured layer-wise sparsity ratios for model pruning. For each task, we apply a dedicated mapping function, to transform the raw IF scores into task-specific layer importance measures. On the contrary, pre-existing methods only use model-only information or heuristics to compute these metrics as indicated by the red arrows.
@@ 3.1 Preliminaries and Notation
$I_{\theta^{*}}(z_{k})\coloneqq\frac{d\theta^{(k)}}{d\epsilon}\bigg{|}_{\epsilon=0}$ $=H(\theta^{*})^{-1}\nabla_{\theta}\ell(z_{k},\theta),$
@@
$I(z_{i})=\sum_{j=1}^{m}\nabla_{\theta}\ell(z^{\mathcal{V}}_{j},\theta)^{\top}H(\theta)^{-1}\nabla_{\theta}\ell(z_{i},\theta).$
@@ 3.2 LayerIF: Estimating Layer Quality via Influence Functions
$I^{(l)}(z_{i})=\sum_{j=1}^{m}\nabla_{\theta^{(l)}}\ell(z^{\mathcal{V}}_{j},\theta)^{\top}\left(H^{(l)}(\theta)\right)^{-1}\nabla_{\theta^{(l)}}\ell(z_{i},\theta)$
For each layer $l$, we aggregate the influence scores across the training set, considering only positively influential samples: $S^{(l)}=\sum_{i=1}^{n}\mathbb{I}[I^{(l)}(z_{i})>0]\cdot I^{(l)}(z_{i})$ where $\mathbb{I}[\cdot]$ is the indicator function. This yields a vector $S\in\mathbb{R}^{N}$, where each element $S^{(l)}$ captures the cumulative positive influence of training data on validation performance through layer $l$.
@@ 3.3 Layer-wise Expert Allocation using LayerIF
We now discuss how we can use our LayerIF method to calculate the optimal distribution of a fixed set of $N$ parallel expert modules. To inform expert allocation, we evaluate the quality of each layer using LayerIF, which measures the sensitivity of model predictions to individual training samples. We compute the IF matrix for each of the Transformer layers in an LLM on a train and validation set. To obtain a scalar score $v_{i}$ for each layer, we aggregate its corresponding IF values while excluding any training samples with negative influence, i.e., those that increase the model’s validation loss, from the final sum.
To map LayerIF scores to expert allocations, we first invert the raw scores obtained for each layer. Since our IF values are inverted in reflecting loss sensitivity, we apply a sign inversion to ensure that layers with higher (i.e., less negative) IF scores remain lower in magnitude after transformation. Formally, for each layer $i$, we compute $\tilde{v}_{i}=-v_{i}$. Next, we apply a power transformation to modulate the sharpness of the allocation. Specifically, we raise the inverted values to a hyperparameter-controlled exponent $\beta>0$, yielding transformed scores $\hat{v}_{i}=\tilde{v}_{i}^{\beta}$. This exponent controls the dispersion of the allocation: higher values of $\beta$ amplify disparities between layers. We then scale the transformed values such that the total allocation, excluding a baseline of one expert per layer, matches the desired target sum $T$. Specifically, we compute $f_{i}=\frac{\hat{v}_{i}}{\sum_{j=1}^{m}\hat{v}_{j}}\cdot(T-m),$ where $m$ is the number of layers, and the floor of each fractional allocation $f_{i}$ is taken with a minimum of 1 added: $s_{i}=\lfloor f_{i}\rfloor+1$. This guarantees that each layer receives at least one expert. Due to flooring, the sum $\sum_{i}s_{i}$ may be less than $T$. We compute the remaining allocation budget $r=T-\sum_{i}s_{i}$ and distribute these remaining units to the $r$ layers with the largest fractional remainders $f_{i}-\lfloor f_{i}\rfloor$, ensuring the total sum constraint is satisfied: $s_{i}\leftarrow s_{i}+1,\quad\text{for top }r\text{ layers with highest }(f_{i}-\lfloor f_{i}\rfloor).$ This procedure produces a final expert allocation vector $S=[s_{1},s_{2},\ldots,s_{m}]$ such that $\sum_{i}s_{i}=T$ and $s_{i}\geq 1$ for all $i$.
@@
A router $S^{i,t}_{j}$, parameterized by a trainable weight matrix $W^{i,t}_{r}\in\mathbb{R}^{n\times N_{j}}$, dynamically assigns experts for a given input $x$. Following the MoLA framework [45], we adopt a top-$K$ routing strategy, where only the $K$ most relevant experts are selected for computation. To encourage balanced expert utilization, we additionally incorporate a load balancing loss at each layer [89]. The routing process is formally defined as $S^{i,t}_{j}(x)=\frac{\text{TopK}(\text{Softmax}(W^{i,t}_{r}x),K)_{j}}{\sum_{j=1}^{K}\text{TopK}(\text{Softmax}(W^{i,t}_{r}x),K)_{j}},$ where $S^{i,t}_{j}(x)$ denotes the normalized assignment weight for expert $j$. The final layer output $h^{i,t}$ is computed as: $h^{i,t}=W^{i,t}_{0}x+\sum_{j=1}^{K}S^{i,t}_{j}(x)A^{i,t}_{j}B^{i,t}_{j}x,$ where $W^{i,t}_{0}$ is a pre-trained base weight matrix, and $A^{i,t}_{j}$, $B^{i,t}_{j}$ are low-rank matrices parameterizing the $j$-th expert. Thus, the output $h^{i,t}$ combines the base transformation of $x$ with the aggregated contributions of the top-$K$ experts, each scaled by its corresponding assignment probability $S^{i,t}_{j}(x)$.
@@ 3.4 Layer Sparsity Allocation for Model Pruning via LayerIF
To ensure stable and meaningful sparsity allocation, we post-process the raw layer-wise IF scores before applying the mapping function. First, we take the absolute value of each score, and normalize these values to the range $[0,1]$ using min-max normalization. This normalization step standardizes the range of IF magnitudes across layers. Next, we apply the Savitzky–Golay filter [91] to the normalized vector $\tilde{s}=(\tilde{s}_{1},\tilde{s}_{2},\dots,\tilde{s}_{L})$ to smooth local fluctuations and reduce noise while preserving the overall trend in layer quality. The smoothed scores are then used as input to the sparsity allocation function $\phi$ described below. This two-step normalization and smoothing process improves robustness and consistency in sparsity assignment, particularly in cases where the raw influence scores vary erratically between adjacent layers.
@@ 4 Experimental Results
Datasets and Base LLMs. For the layer-wise expert allocation in the LoRA-MoE application, we assess post-training zero-shot performance on two GLUE datasets and three commonsense reasoning tasks, following a protocol similar to that of Qing et al. [46]. The datasets used include: MRPC [93], CoLA [94], ScienceQA [95], CommonsenseQA [96], and OpenBookQA [97]. For layer sparsity allocation for model pruning we again assess the post-pruning zero-shot performance of our models on several NLP tasks. Namely, BoolQ [98], Hellaswag [99], Winogrande [100], ARC Easy and ARC Challenge [101], RTE [94], and OpenBookQA [97]. For both of the applications we conduct experiments on Mistral-7b-v0.3 [7] and Gemma-7b [8]. All of our experiments run on 8$\times$×NVIDIA RTX 6000 Ada GPUs. All code and reproducibility configurations are provided in Appendix G and we provide further information on statistical significance of our results in F.
@@
Additionally, we compare against AlphaLora [46], and 3 baselines from MoLA [45]: MoLA-2468, MoLA-$5$ and MoLA-8642.
4.1 Results on Layer-wise Expert Allocation using LayerIF
@@
All LayerIF configurations outperform baselines. Table 1 presents accuracy across datasets for expert allocation methods, including AlphaLora, three variants of MoLA, and our proposed LayerIF configurations. LayerIF consistently outperforms prior approaches, achieving the highest average accuracy across datasets. Specifically, the best performing LayerIF configuration had a relative percentage increase of 1.61% over the best performing baseline and the average of the LayerIF configurations had a 3.54% percentage increase over the average of other strong baselines. These results demonstrate that IF-based layer quality estimation provides a more effective signal for expert allocation in LLMs.
Consistently strong performance across varying number of experts. In Figure 2, we provide average results for varying the number of experts, specifically 80, 160 and 224 experts on Mistral-7b-v0.1. MoLA(1234) and MoLA(46810) are introduced as additional baselines, following the most performant pattern as claimed in the MoLA paper. We can clearly see that LayerIF clearly outperforms the baselines in all three settings by having an average accuracy that is a 3.64% percent increase over the second best average accuracy AlphaLora. Expert allocation results for Gemma-7b are provided in Appendix C.
@@ 4.2 Results on Layer Sparsity Allocation for Model Pruning via LayerIF
Robust performance across multiple pruning ratios. Figure 3 demonstrates the consistently reliable performance of LayerIF across multiple pruning ratios. LayerIF consistently achieved the highest accuracy across all pruning ratios, while baseline methods exhibited greater performance variability, frequently changing ranks across sparsity levels. On average, LayerIF outperformed the next best method, OWL, by a 0.37% percentage increase. Moreover, we provide additional layer-wise sparsity allocation results for Gemma-7b in Appendix D and results with and without the smoothing of IF values in Appendix E.
5 Discussion
@@
Computational Complexity. One limitation of IF-based methods in general (at test-time) is their high computational complexity, primarily due to the Hessian inversion. However, optimizing this compute time is an active area of research [28, 82] and our LayerIF framework can employ any efficient IF method proposed in the future. Additionally, it is important to note that our framework is not for use during inference. That is, since our framework is applied during model training, the computational overhead of IF is a minuscule fraction of the time it takes to pretrain LLMs.
@@
Comparison with Hessian-free IF Methods. While Hessian-free IF methods are computationally more efficient, they generally lead to a drop in performance [37, 28]. To assess if Hessian-free methods such as TracIn [27] can be used for layer quality influence estimation, we compute the layer-wise expert allocations for each dataset from our main experiments using TracIn as the IF method in LayerIF and compare these allocations with those derived from the Hessian-based method, DataInf [81] (employed in our main experiments). We compute the Spearman’s correlation coefficient and present our findings in Table 3. We find that Hessian-free methods fail to capture meaningful information about layer-wise training quality as there is little to no agreement in the resulting layer allocations across datasets when employing TracIn compared to DataInf in LayerIF.

### v3

@@ 1 Introduction
There is a broad effort in the ML community to interpret and explain the behavior of these complex models. Many approaches take a model-centric perspective, examining how the model’s learned weights effect predictions [20, 19]. Other methods use activations, or attention patterns to deduce internal mechanisms [21, 22, 23, 24]. In contrast Influence Functions (IFs) [25, 26] offer a more explicitly data-centric perspective by quantifying the effect of individual training points on model’s loss landscape. Rather than focusing on internal representations, Influence Functions (IFs) measure how a model’s output would change if a given training example was perturbed or removed, effectively tracing the influence of training data on the model’s predictions. They have proven effective in a range of downstream machine learning tasks, including mislabeled data detection [27, 28, 29], optimal subset selection [30, 31, 32], model interpretation [33, 34, 35], data attribution [36], data valuation [37], in-context learning [38], and analyzing model biases [39, 40].
@@
Figure 1: We present an overview of the LayerIF pipeline via the green arrows. We quantify per-layer quality in a pretrained LLM using Influence Functions (IFs), demonstrating that the same model produces distinct layer-wise IF scores across different datasets, revealing dataset-specific specialization. The pipeline begins by extracting gradients at each Transformer block of the LLM, computed separately for each dataset. These gradients and the dataset are then used to estimate layer-wise IF scores that serve as data-driven proxies for layer quality. To demonstrate the utility of these scores, we consider two downstream applications: (a) the allocation of optimal experts per layer in a LoRA-MoE architecture, and (b) the computation of structured layer-wise sparsity ratios for model pruning. For each task, we apply a dedicated mapping function, to transform the raw IF scores into task-specific layer importance measures. On the contrary, pre-existing methods only use model-only information or heuristics to compute these metrics as indicated by the red arrows. Figure 1 is a conceptual illustration contrasting our method with the baselines; the shown values are illustrative, not experimental.
@@ 3.1 Preliminaries and Notation
$I_{\theta^{*}}(z_{k})\coloneqq\frac{d\theta^{(k)}}{d\epsilon}\bigg|_{\epsilon=0}$ $=-H(\theta^{*})^{-1}\nabla_{\theta}\ell(z_{k},\theta),$
@@
$I(z_{i})=-\sum_{j=1}^{m}\nabla_{\theta}\ell(z^{\mathcal{V}}_{j},\theta)^{\top}H(\theta)^{-1}\nabla_{\theta}\ell(z_{i},\theta).$
@@ 3.2 LayerIF: Estimating Layer Quality via Influence Functions
$I^{(l)}(z_{i})=-\sum_{j=1}^{m}\nabla_{\theta^{(l)}}\ell(z^{\mathcal{V}}_{j},\theta)^{\top}\left(H^{(l)}(\theta)\right)^{-1}\nabla_{\theta^{(l)}}\ell(z_{i},\theta)$
For each layer $l$, we aggregate the influence scores across the training set, considering only positively influential samples after a sign inversion  [^1] This sign inversion is done since it is more intuitive that positive influential samples are beneficial to the model.: $S^{(l)}=\sum_{i=1}^{n}\mathbb{I}[I^{(l)}(z_{i})>0]\cdot I^{(l)}(z_{i})$ where $\mathbb{I}[\cdot]$ is the indicator function. This yields a vector $S\in\mathbb{R}^{N}$, where each element $S^{(l)}$ captures the cumulative positive influence of training data on validation performance through layer $l$.
@@ 3.3 Layer-wise Expert Allocation using LayerIF
We now discuss how we can use our LayerIF method to calculate the optimal distribution of a fixed set of $N$ parallel expert modules. To inform expert allocation, we evaluate the quality of each layer using LayerIF, which measures the sensitivity of model predictions to individual training samples. We compute the IF matrix for each of the Transformer layers in an LLM on a train and validation set. To obtain a scalar score $v_{i}$ for each layer, we aggregate its corresponding IF values while excluding any training samples with negative influence, i.e., those that decrease the model’s loss, from the final sum.
To map LayerIF scores to expert allocations, we first invert the raw scores obtained for each layer. Since our IF values are inverted to reflect loss sensitivity, we apply a sign inversion to ensure that layers with higher (i.e., less negative) IF scores remain lower in magnitude after transformation. Formally, for each layer $i$, we compute $\tilde{v}_{i}=-v_{i}$ where $v_{i}=-S^{(l)}$. Next, we apply a power transformation to modulate the sharpness of the allocation. Specifically, we raise the inverted values to a hyperparameter-controlled exponent $\beta>0$, yielding transformed scores $\hat{v}_{i}=\tilde{v}_{i}^{\beta}$. This exponent controls the dispersion of the allocation: higher values of $\beta$ amplify disparities between layers. We then scale the transformed values such that the total allocation, excluding a baseline of one expert per layer, matches the desired target sum $T$. Specifically, we compute $f_{i}=\frac{\hat{v}_{i}}{\sum_{j=1}^{m}\hat{v}_{j}}\cdot(T-m),$ where $m$ is the number of layers, and the floor of each fractional allocation $f_{i}$ is taken with a minimum of 1 added: $s_{i}=\lfloor f_{i}\rfloor+1$. This guarantees that each layer receives at least one expert. Due to flooring, the sum $\sum_{i}s_{i}$ may be less than $T$. We compute the remaining allocation budget $r=T-\sum_{i}s_{i}$ and distribute these remaining units to the $r$ layers with the largest fractional remainders $f_{i}-\lfloor f_{i}\rfloor$, ensuring the total sum constraint is satisfied: $s_{i}\leftarrow s_{i}+1,\quad\text{for top }r\text{ layers with highest }(f_{i}-\lfloor f_{i}\rfloor).$ This procedure produces a final expert allocation vector $S=[s_{1},s_{2},\ldots,s_{m}]$ such that $\sum_{i}s_{i}=T$ and $s_{i}\geq 1$ for all $i$.
@@
A router $S^{i,t}_{j}$, parameterized by a trainable weight matrix $W^{i,t}_{r}\in\mathbb{R}^{n\times N_{j}}$, dynamically assigns experts for a given input $x$. Following the MoLA framework [46], we adopt a top-$K$ routing strategy, where only the $K$ most relevant experts are selected for computation. To encourage balanced expert utilization, we additionally incorporate a load balancing loss at each layer [89]. The routing process is formally defined as $S^{i,t}_{j}(x)=\frac{\text{TopK}(\text{Softmax}(W^{i,t}_{r}x),K)_{j}}{\sum_{j=1}^{K}\text{TopK}(\text{Softmax}(W^{i,t}_{r}x),K)_{j}},$ where $S^{i,t}_{j}(x)$ denotes the normalized assignment weight for expert $j$. The final layer output $h^{i,t}$ is computed as: $h^{i,t}=W^{i,t}_{0}x+\sum_{j=1}^{K}S^{i,t}_{j}(x)A^{i,t}_{j}B^{i,t}_{j}x,$ where $W^{i,t}_{0}$ is a pre-trained base weight matrix, and $A^{i,t}_{j}$, $B^{i,t}_{j}$ are low-rank matrices parameterizing the $j$-th expert. Thus, the output $h^{i,t}$ combines the base transformation of $x$ with the aggregated contributions of the top-$K$ experts, each scaled by its corresponding assignment probability $S^{i,t}_{j}(x)$. A step-by-step walkthrough of our algorithm is provided in Appendix C for greater clarity.
@@ 3.4 Layer Sparsity Allocation for Model Pruning via LayerIF
To ensure stable and meaningful sparsity allocation, we post-process the raw layer-wise IF scores before applying the mapping function. First, we take the absolute value of each score, and normalize these values to the range $[0,1]$ using min-max normalization. This normalization step standardizes the range of IF magnitudes across layers. Next, we apply the Savitzky–Golay filter [91] to the normalized vector $\tilde{s}=(\tilde{s}_{1},\tilde{s}_{2},\dots,\tilde{s}_{L})$ to smooth local fluctuations and reduce noise while preserving the overall trend in layer quality  [^2] Raw Influence Function (IF) scores at adjacent layers exhibit high-frequency noise due to stochastic Hessian approximations and mini-batch gradient estimates. This variance is well documented by [34], whose layer–token heatmaps show similar oscillations.. The smoothed scores are then used as input to the sparsity allocation function $\phi$ described below. This two-step normalization and smoothing process improves robustness and consistency in sparsity assignment, particularly in cases where the raw influence scores vary erratically between adjacent layers.
@@ 4 Experimental Results
[…]
@@
Additionally, we compare against AlphaLora [47], and 3 baselines from MoLA [46]: MoLA-2468, MoLA-5555 and MoLA-8642.
@@
4.1 Results on Layer-wise Expert Allocation using LayerIF
All LayerIF configurations outperform baselines. Table 1 presents accuracy across datasets for expert allocation methods, including AlphaLora, three variants of MoLA, and our proposed LayerIF configurations. LayerIF consistently outperforms prior approaches, achieving the highest average accuracy across datasets. Specifically, the best performing LayerIF configuration had a relative percentage increase of 1.61% over the best performing baseline and the average of the LayerIF configurations had a 3.52% percentage increase over the average of other strong baselines. These results demonstrate that IF-based layer quality estimation provides a more effective signal for expert allocation in LLMs.
Consistently strong performance across varying number of experts. In Figure 2, we provide average results for varying the number of experts, specifically 80, 160 and 224 experts on Mistral-7b-v0.1. MoLA(1234) and MoLA(46810) are introduced as additional baselines, following the most performant pattern as claimed in the MoLA paper. We can clearly see that LayerIF clearly outperforms the baselines in all three settings by having an average accuracy that is a 3.64% percent increase over the second best average accuracy AlphaLora. Expert allocation results for Gemma-7b are provided in Appendix D.
@@ 4.2 Results on Layer Sparsity Allocation for Model Pruning via LayerIF
Robust performance across multiple pruning ratios. Figure 3 demonstrates the consistently reliable performance of LayerIF across multiple pruning ratios. LayerIF consistently achieved the highest accuracy across all pruning ratios, while baseline methods exhibited greater performance variability, frequently changing ranks across sparsity levels. On average, LayerIF outperformed the next best method, OWL, by a 0.37% percentage increase. Moreover, we provide additional layer-wise sparsity allocation results for Gemma-7b in Appendix E and results with and without the smoothing of IF values in Appendix F.
5 Discussion
Interpretation of Layer Sensitivity. Influence Functions quantify how small perturbations in training data affect the model’s loss. In this view, layers that are well-trained and have effectively converged tend to exhibit lower sensitivity to such perturbations, as reflected by smaller IF magnitudes. To further examine this relationship, we conduct an ablation in which we invert our allocation strategy, assigning higher pruning ratios to layers with lower IF sensitivity. Under a 70% sparsity budget on Mistral-7b-v0.1, this reversed allocation leads to consistently worse downstream performance, reinforcing that layers with lower IF sensitivity are better trained and should be preserved more aggressively, as can be seen in Table 3.
Table 3: Comparison of accuracies across different pruning strategies (Magnitude, Wanda, SparseGPT) between LayerIF and reversed allocation at 70% pruning on Mistral-7b-v0.1.
Method Magnitude Wanda SparseGPT Average
LayerIF 33.44 32.50 41.14 35.69
Reversed Allocation 33.10 32.02 38.45 34.52
Guidelines for Influence Sample Selection. We provide a empirical perspective on how to select the subset of influence samples used to estimate layer quality. Throughout the paper, we aggregate only positive-influence samples (+VE) by default, as they reflect training points that reduce validation loss and thus carry constructive learning signal. In Section 4, we additionally examine two variants: (ALL) which includes all samples regardless of sign, and (TOP-25%) which retains the top quartile of positive scores by magnitude. Empirically, the TOP-25% variant achieves the highest average performance across datasets, suggesting it is a strong default in practice.
From a theoretical standpoint, this selection serves as a noise-filtering mechanism. Because IF estimates are computed from stochastic Hessian and gradient approximations, small-magnitude scores often correspond to spurious or weak training–validation interactions. Retaining only the upper tail of positive influences increases the signal-to-noise ratio. Prior studies have also shown that IF values in deep models are heavy-tailed [34], implying that a small fraction of training samples account for most of the total positive influence. Hence, focusing on this high-influence subset isolates the most informative training signals.
In practice, one can determine the cutoff ratio using a simple cumulative influence curve: sorting positive IF values in descending order and identifying the elbow point where marginal contribution drops sharply provides a data-driven threshold. This adaptive thresholding remains an exciting direction for future work.
@@ 5 Discussion
Computational Complexity. One limitation of IF-based methods in general (at test-time) is their high computational complexity, primarily due to the Hessian inversion. However, optimizing this compute time is an active area of research [29, 82] and our LayerIF framework can employ any efficient IF method proposed in the future. Additionally, it is important to note that our framework is not for use during inference. That is, since our framework is applied during model training, the computational overhead of IF is a minuscule fraction of the time it takes to pretrain LLMs. Detailed wall-clock times and memory consumption for varying model and dataset sizes are presented in Appendix H.
@@
Comparison with alternative IF Methods. One class of alternative IF methods are Hessian-free methods like TracIn [28]. While Hessian-free IF methods are computationally more efficient, they generally lead to a drop in performance [38, 29]. To assess if Hessian-free methods such as TracIn [28] can be used for layer quality influence estimation, we compute the layer-wise expert allocations for each dataset from our main experiments using TracIn as the IF method in LayerIF and compare these allocations with those derived from the Hessian-based method, DataInf [81] (employed in our main experiments). We compute the Spearman’s correlation coefficient and present our findings in Table 4. We find that Hessian-free methods fail to capture meaningful information about layer-wise training quality as there is little to no agreement in the resulting layer allocations across datasets when employing TracIn compared to DataInf in LayerIF. Another class of alternative IF methods are Gauss-Newton Hessians like the EK-FAC method [34]. The main bottleneck with the Gauss-Newton Hessian is the time required for IF computation. In our experiments, the Gauss Newton Hessian was roughly 160× slower than using the DataInf approximation, making DataInf the only practical choice for our large-scale experiments.